Stock A's stock has a beta of 1.20, and its required return is 12.25%. Stock B's beta is 0.80. If the risk-free rate is 4.25%, what is the required rate of return on B's stock?

8.76%

9.21%

10.25%

9.58%

8.12%
Answer #1

Required rate=risk free rate+Beta*market risk premium

A:

12.25=4.25+1.2*market risk premium

market risk premium=(12.25-4.25)/1.2

=6.67%(Approx).

B:

Required rate=4.25+(0.8*6.67)

=9.58%(Approx).